how do you find the slant hight of a cone with the altitiude of 12 and radias of 9??
Answer by checkley75(3666)

9^2+12^2=S^2
81+144=S^2
225=S^2
S=SQRT225
S=15 ANS. FOR THE SLANT HEIGHT.